13sqrt3 + 4sqrt150 - 7sqrt48 + 2 sqrt2. Please help simplify. I got it down to 24sqrt6 + 13sqrt3 -14sqrt6sqrt2. Not sure if you can combine numbers if there are two square roots next to each other. For example, sqrt6 and then sqrt6sqrt2. Can those 6's be combined? Thanks so much!
Answer by richwmiller(17219)

2sqrt(2)-15sqrt(3)+20sqrt(6)=25.837
We start with:
13sqrt3 + 4sqrt150 - 7sqrt48 + 2 sqrt2
We can reduce 4sqrt(150) and -7sqrt(48)
4sqrt(150)=4sqrt(25*6)=20sqrt(6)
-7sqrt(48)=-7sqrt(16*3)=-7*(4)sqrt(3)
13sqrt(3)-28sqrt(3)=-15sqrt(3)
We end up with
2sqrt(2)-15sqrt(3)+20sqrt(6)=25.837
I have no idea what you did.